# Reinforcement Learning Pair Trading: A Dynamic Scaling Approach

### v1

@@ 5.2.1 Result Comparison
Table 5: Evaluation Metrics Comparison between Trading Techniques
Gatev et al. (2006) Kim & Kim (2019) RL1 RL2
@@
Cumulative Return 8.33% -0.16% -35.16% -35.79% 1.89% 9.94% -31.99% -77.81% 31.53% -87.12%
CAGR 195.12% -2.19% -99.71% -99.75% 30.05% 278.72% -99.56% -100.00% 3974.65% -100.00%
Sharpe Ratio 25.91 -1.67 -2.04 -2.60 5.44 32.74 -8.77 -1.99 94.34 -1.93
@@
Max Win Action (\$) 75.35 163.52 606.75 606.75 43.72 121.74 70.59 307.78 648.87 160.15
Max Loss Action (\$) -27.73 -187.86 -763.70 -553.25 -108.51 -21.33 -282.84 -389.22 -64.97 -1456.43
Avg Win Action P&L (\$) 14.50 41.72 38.68 37.47 5.51 17.77 11.30 15.88 90.94 8.03
Avg Loss Action P&L (\$) -2.90 -54.68 -58.75 -60.78 -3.28 -7.06 -24.95 -17.78 -21.00 -23.49
@@
Skew 1840 -54 -358 -358 -874 2673 -1899 -374 4314 -3048
Kurtosis 48145 135 4138 4201 133603 114944 54808 12987 254283 138851
  • 1
    The trading period is from 01-Dec-2023 to 31-Dec-2023.
  • 2
    The transaction cost is 0.02%, and the interest rate is 5.5%.
  • 3
    RL 1 stands for the pair trading that allows Reinforcement Learning to decide upon the investment timing.
  • 4
    RL 2 stands for Reinforcement Learning pair trading that allows the RL agent to decide both investment timing and quantity.
(a) Gatev et al. (2006) Pair Trading
(b) Kim & Kim (2019) Pair Trading (PPO)
(c) RL 1 Pair Trading (A2C)
(d) RL 2 Pair Trading (A2C)
Figure 6: The portfolio value trend comparison between agents with the best performance
Our results are presented in Table 5. The results display a positive return for traditional pair trading technique Gatev et al. [2006]. The algorithm A2C displays a positive return for RL pair trading techniques. However, algorithms PPO, SAC, and DQN do not perform as well as A2C. If we view A2C as the chosen algorithm for pair trading, the results show a steady income from pair trading. The traditional pair trading of Gatev et.al. 2006 displays a stable income compared to others due to its rule-based execution stability.
The first adoption of RL 1 pair trading is close to the traditional method. The result table shows that it achieved much better results than the traditional pair trading approach Gatev et al. [2006]. The second adoption of RL 2 pair trading is significantly different from the RL 1 trading that decides only timing, which produces more profit than other techniques under the same level of volatility. Kim & Kim (2019)’s method did not achieve a positive return. Since the method was developed for the forex market, it has not adapted well to the extremely volatile cryptocurrency world.
Behaviour-wise, PPO, DQN and SAC tend to conduct excessive transactions that are not profitable. On the contrary, A2C have fewer trades but higher profits on each trade. The RL 2 pair trading shows further fewer total actions because of the adjust position action, where we do not consider a position adjustment as one trade until it is closed. Apart from the result in Table 5, the portfolio growth trend with the best performing RL algorithm agent is presented in Figure 6. Most of the pair trading experiments, including Gatev et al. [2006], RL1, and RL2, display a stable upturn, which is ideal from the perspective of pair trading. We can see from the drawdown graphs that RL 1 produces fewer drawdowns compared to the non-RL pair trading method from Gatev et al. [2006], with even more profit. RL2 displays the strongest profitability as well as the largest drawdown. In general, all three pair trading methods show the ability to generate stable income in a volatile trading market.
@@ 5.2.2 Effect of Transaction Cost
The profitability of high-frequency trading techniques always receives a significant impact from transaction costs. Cryptocurrency exchanges normally provide a large-volume discount scheme. The Binance fee ranges from 0.02% to even 0%, depending on the volume and the holding of their membership token. Considering that the users of these techniques may benefit from different transaction fee tiers, we explore the trading techniques under different transaction cost tiers as well.
@@
Max Win Action (\$) 72.82 114.76 43.36 606.22
Max Loss Action (\$) -30.26 -169.52 -8.30 -168.81
Avg Win Action P&L (\$) 13.57 37.94 16.10 70.99
Avg Loss Action P&L (\$) -4.99 -47.68 -5.64 -48.28
0.01% Transaction Fee Gatev et al. (2006) Kim & Kim (2019) RL1 RL2
Cumulative Profit 9.43% -1.13% 9.88% 33.99%
Sharpe Ratio 29.84 -7.07 33.24 104.40
@@
Max Win Action (\$) 76.20 65.99 121.74 675.23
Max Loss Action (\$) -26.88 -169.66 -21.33 -27.91
Avg Win Action P&L (\$) 13.93 24.88 17.52 98.74
Avg Loss Action P&L (\$) -2.48 -44.81 -7.06 -10.79
0% Transaction Fee Gatev et al. (2006) Kim & Kim (2019) RL1 RL2
Cumulative Profit 10.54% -2.00% 9.94% 80.92%
Sharpe Ratio 33.90 -5.76 32.74 2668.86
@@
Max Win Action (\$) 77.04 163.59 121.74 699.51
Max Loss Action (\$) -26.03 -217.54 -21.33 -72.21
Avg Win Action P&L (\$) 13.07 36.69 17.77 104.25
Avg Loss Action P&L (\$) -2.43 -80.76 -7.06 -16.68
Table 6: Evaluation Metrics Comparison under Different Transaction Cost
  • 1
    The trading period is from 01-Dec-2023 to 31-Dec-2023 with interest rate as 5.5%.
  • 2
    RL 1 stands for the pair trading that allows Reinforcement Learning to decide upon the investment timing.
  • 3
@@
    RL 2 stands for Dynamic Scaling Reinforcement Learning pair trading that allows the RL agent to decide both investment timing and quantity.
  • 4
    We adopted algorithm PPO for Kim & Kim (2019), and A2C for RL1 and RL2.
With more exploration under 0.05%, 0.01% and 0% transaction costs compared to the default 0.02% transaction cost, we can see the significant impact of decreasing the transaction cost. The participating approaches are adopted with the most profitable algorithm based on the backtesting result (Table 5). We can see that trading techniques generally perform better under lower transaction costs. RL-based techniques tend to perform more trades when the transaction costs are lower.
6 Discussion and conclusions
Pair trading has been a popular algorithmic trading method for decades. The in-demand high-frequency trading domain requires a fast-track decision-making process. However, the traditional rule-based pair trading technique lacks the flexibility to cater to volatile market movement. In this research, we proposed a mechanism to adopt Reinforcement Learning (RL)to observe the market and produce profitable pair trading decisions. The the first adoption of Reinforcement Learning pair trading grants the RL 1 agent the flexibility to decide action timing. The second adoption of Reinforcement Learning 2 Pair Trading further gives RL agent the access to decide the timing and invest quantity.
We compared it to the traditional rule-based Pair Trading technique [Gatev et al., 2006] and a state-of-the-art RL pair trading technique [Kim & Kim, 2019], for 2023-Dec in the cryptocurrency market for BTCEUR and BTCGBP under standard future 0.02% transaction cost. Kim & Kim’s method does not perform well in the cryptocurrency world. The Gatev et al.’s method achieved 8.33% per trading period. our first adoption of RL 1 method achieved 9.94%, and the second adoption of RL 2 method achieved 31.53% returns during the trading period. The outperformance is generally consistent across different transaction costs. The evaluation metrics show that RL-based techniques are generally more active than traditional techniques in the cryptocurrency market under various transaction costs. In general, our trading methods have greater market participation than Gatev et al.’s traditional rule-based pair trading and Kim & Kim’s threshold-adaptive RL Pair Trading (Table 56).
@@ 6 Discussion and conclusions
Comparison between RL-based pair trading revealed the relationship between profitability and actions. Because financial trading is a special case of RL environment, every action in financial trading is punished by the transaction cost. We notice that profitable RL trading often has less total trade count and higher profit per win trade. That means the RL is better at spotting chances to make higher profits. The RL 2 pair trading produces higher profits because of higher average wins from the position adjustment mechanism.
The techniques open some future work opportunities. We can develop the RL pair trading with applications to multi-leg strategies. At this moment, pair trading either relies on forming a pool of assets into a two-leg pair and executing trades upon the selected pair or relies on an optimisation-based analytical function for pair combining them amongst a selected pool. We believe the RL optimiser might provide better execution of pair formation. The breakdown of Pair Formation from preliminary preparation into part of trading could significantly boost the trading efficiency, diversify the asset holding risk and potentially enhance profitability.
Acknowledgements

### v2

@@ 5.2.1 Result Comparison
Our results are presented in Table 5. The results display a positive return for the traditional pair trading technique Gatev et al. (2006). The algorithm A2C displays a positive return for RL pair trading techniques. However, the PPO, SAC, and DQN algorithms do not perform as well as A2C. If we view A2C as the chosen algorithm for pair trading, the results show a steady income from pair trading. The traditional pair trading of Gatev et al. (2006) displays a stable income compared to others due to its rule-based execution stability.
[…]
Table 5: Evaluation metrics comparison between trading techniques.
Gatev et al. (2006) Kim and Kim (2019) RL1 RL2
@@
Cumulative Return 8.33% --0.16% --35.16% --35.79% 1.89% 9.94% --31.99% --77.81% 31.53% --87.12%
CAGR 195.12% --2.19% --99.71% --99.75% 30.05% 278.72% --99.56% --100.00% 3974.65% --100.00%
Sharpe Ratio 25.91 --1.67 --2.04 --2.60 5.44 32.74 --8.77 --1.99 94.34 --1.93
@@
Max Win Action (USD) 75.35 163.52 606.75 606.75 43.72 121.74 70.59 307.78 648.87 160.15
Max Loss Action (USD) --27.73 --187.86 --763.70 --553.25 --108.51 --21.33 --282.84 --389.22 --64.97 --1456.43
Avg Win Action Profit/Loss (USD) 14.50 41.72 38.68 37.47 5.51 17.77 11.30 15.88 90.94 8.03
Avg Loss Action Profit/Loss (USD) --2.90 --54.68 --58.75 --60.78 --3.28 --7.06 --24.95 --17.78 --21.00 --23.49
@@
Skew 1840 --54 --358 --358 --874 2673 --1899 --374 4314 --3048
Kurtosis 48,145 135 4138 4201 133,603 114,944 54,808 12,987 254,283 138,851
{adjustwidth}
[…]
Behavior-wise, PPO, DQN, and SAC tend to conduct excessive transactions that are not profitable. On the contrary, A2C have fewer trades but higher profits on each trade. RL 2 pair trading shows further fewer total actions because of the adjusted position action, where we do not consider a position adjustment as one trade until it is closed. Apart from the result in Table 5, the portfolio growth trend with the best-performing RL algorithm agent is presented in Figure 6 (a comparison with Gatev et al. (2006) is provided in Figure 7a in the Appendix). Most of the pair trading experiments, including Gatev et al. (2006), RL1, and RL2, display a stable upturn, which is ideal from the perspective of pair trading. From the drawdown graphs, we can observe that RL 1 produces fewer drawdowns compared to the non-RL pair trading method from Gatev et al. (2006) and has a significantly higher win/loss action ratio due to differences in threshold settings. However, RL1’s cumulative profit is not consistently higher, and when transaction fees are zero, its cumulative profit is slightly lower than that of the Gatev et al. method. RL2 displays the strongest profitability, despite a lower win/loss action ratio, due to its progressive trading strategy. In general, all three pair trading methods show the ability to generate stable income in a volatile trading market.
a RL 1 Pair Trading (A2C).
b RL 2 Pair Trading (A2C).
Figure 6: Comparison of portfolio value trends for RL 1 and RL 2 Pair Trading (A2C).
@@ 5.2.2 Effect of Transaction Cost
The profitability of high-frequency trading techniques is always significantly impacted by transaction costs. Cryptocurrency exchanges normally provide a large-volume discount scheme. The Binance fee ranges from 0.02% to even 0%, depending on the volume and the holding of their membership token. Considering that the users of these techniques may benefit from different transaction fee tiers, we explore the trading techniques under different transaction cost tiers as well. With more exploration under 0.05%, 0.01%, and 0% transaction costs compared to the default 0.02% transaction cost, we can see the significant impact of decreasing the transaction cost. The participating approaches are adopted with the most profitable algorithm based on the backtesting result (Table 5). We can see that trading techniques generally perform better under lower transaction costs. RL-based techniques tend to perform more trades when the transaction costs are lower.
6 Discussion and Conclusions
Pair trading has been a popular algorithmic trading method for decades. The in-demand high-frequency trading domain requires a fast-track decision-making process. However, the traditional rule-based pair trading technique lacks the flexibility to cater to volatile market movement. In this research, we proposed a mechanism to adopt Reinforcement Learning (RL)to observe the market and produce profitable pair trading decisions. The first adoption of Reinforcement Learning pair trading grants the RL 1 agent the flexibility to decide action timing. The second adoption of Reinforcement Learning 2 pair trading further gives the RL agent the access to decide the timing and invest quantity.
We compared it to the traditional rule-based pair trading technique (Gatev et al., 2006) and a state-of-the-art RL pair trading technique (Kim and Kim, 2019) for December 2023 in the cryptocurrency market for BTCEUR and BTCGBP under a standard future 0.02% transaction cost. Kim and Kim’s method does not perform well in the cryptocurrency world. Gatev et al.’s method achieved 8.33% per trading period. Our first adoption of the RL 1 method achieved 9.94%, and the second adoption of the RL 2 method achieved 31.53% returns during the trading period. The outperformance is generally consistent across different transaction costs. The evaluation metrics show that RL-based techniques are generally more active than traditional techniques in the cryptocurrency market under various transaction costs. In general, our trading methods have greater market participation than Gatev et al.’s traditional rule-based pair trading and Kim and Kim’s threshold-adaptive RL pair trading (Tables 5 and 6).
Table 6: Evaluation metrics comparison under different transaction costs.
@@ 6 Discussion and Conclusions
Max Win Action (USD) 72.82 114.76 43.36 606.22
Max Loss Action (USD) --30.26 --169.52 --8.30 --168.81
Avg Win Action Profit/Loss (USD) 13.57 37.94 16.10 70.99
Avg Loss Action Profit/Loss (USD) --4.99 --47.68 --5.64 --48.28
0.01% Transaction Fee Gatev et al. (2006) Kim and Kim (2019) RL1 RL2
Cumulative Profit 9.43% --1.13% 9.88% 33.99%
Sharpe Ratio 29.84 --7.07 33.24 104.40
@@
Max Win Action (USD) 76.20 65.99 121.74 675.23
Max Loss Action (USD) --26.88 --169.66 --21.33 --27.91
Avg Win Action Profit/Loss (USD) 13.93 24.88 17.52 98.74
Avg Loss Action Profit/Loss (USD) --2.48 --44.81 --7.06 --10.79
0% Transaction Fee Gatev et al. (2006) Kim and Kim (2019) RL1 RL2
Cumulative Profit 10.54% --2.00% 9.94% 80.92%
Sharpe Ratio 33.90 --5.76 32.74 2668.86
@@
Max Win Action (USD) 77.04 163.59 121.74 699.51
Max Loss Action (USD) --26.03 --217.54 --21.33 --72.21
Avg Win Action Profit/Loss (USD) 13.07 36.69 17.77 104.25
Avg Loss Action Profit/Loss (USD) --2.43 --80.76 --7.06 --16.68
@@
The trading period is from 1 December 2023 to 31 December 2023 with an interest rate of 5.5%. RL 1 stands for the pair trading that allows Reinforcement Learning to decide upon the investment timing. RL 2 stands for dynamic scaling Reinforcement Learning pair trading that allows the RL agent to decide both investment timing and quantity. We adopted the PPO algorithm from Kim and Kim (2019) and A2C for RL1 and RL2.
@@
Comparison between RL-based pair trading revealed the relationship between profitability and actions. Because financial trading is a special case of the RL environment, every action in financial trading is punished by the transaction cost. We notice that profitable RL trading often has a lower total trade count and higher profit per-win trade. That means the RL is better at spotting chances to make higher profits. RL 2 pair trading produces higher profits because of higher average wins from the position adjustment mechanism. When we adopt the righteous trading method, market volatility and transaction cost play crucial roles in profitable trading. Variable thresholds might not be adaptive to highly volatile markets, and fixed-threshold pair trading could lead to missing trading opportunities. RL with dynamic scaling investment could be a good direction in volatile market conditions if low transaction costs are achievable.
The techniques presented have certain limitations and offer opportunities for future work. One limitation is the relatively limited dataset scope, which could be expanded to include more diverse assets and longer timeframes to improve generalization. Additionally, focusing only on two-leg strategies restricts the potential for complex arbitrage opportunities; expanding to multi-leg strategies would enhance robustness. The computational demand during training can also be resource-intensive, requiring system parameter tuning. The model lacks consideration for transaction costs, which might impact real-world profitability. A lack of direct comparison with traditional models is another shortcoming. Future work could involve developing the Reinforcement Learning (RL) approach to multi-leg strategies, integrating pair formation into the trading process, cross-validating across different environments, and experimenting with alternative reward functions to improve decision-making and risk management.
[…]
Conceptualization, A.M. and H.Y.; methodology, A.M. and H.Y.; software, A.M. and H.Y.; validation, H.Y.; formal analysis, H.Y.; investigation, H.Y.; resources, A.M. and H.Y.; data curation, H.Y.; writing—original draft preparation, H.Y.; writing—review and editing, A.M.; visualization, H.Y.; supervision, A.M.; project administration, A.M.; funding acquisition, Not applicable. All authors have read and agreed to the published version of the manuscript.
[…]
This research received no external funding
[…]
Not applicable.
[…]
The original data presented in the study are openly available from Binance Exchange accessed on 8 November 2024 at (https://data.binance.vision/).
[…]
The authors declare no conflicts of interest.
[…]
Appendix A
a
b
Figure 7: Comparison of Pair Trading strategies from (a) Gatev et al. (2006) and (b) Kim and Kim (2019).
{adjustwidth}
[…]